OpenStudy (anonymous):

Find the slope of a line parallel to the line through the given points. M(5, 5), N(6, -2) slope = -7/1 CORRECT?

OpenStudy (anonymous):

@ghazi

hero (hero):

Using y = mx + b 5 = 5m + b -2 = 6m + b b = 5 - 5m b = -2 - 6m b = b 5 - 5m = -2 - 6m 6m - 5m = -2 - 5 m = -7 -7 = -7/1 so correct